OpenStudy (anonymous):

determine the standard form of the equation of the line that passes through (6,0) and (2,-7)

OpenStudy (hba):

\[y-y_1=\frac{ y_2-y_1 }{ x_2-x_1 }(x-x_1)\] Here (x1,y1)=(6,0) (x2,y2)= (2,-7)